# LLM-Resistant Math Word Problem Generation via Adversarial Attacks

### v1

@@ Abstract
Large language models (LLMs) have significantly transformed the educational landscape. As current plagiarism detection tools struggle to keep pace with LLMs’ rapid advancements, the educational community faces the challenge of assessing students’ true problem-solving abilities in the presence of LLMs. In this work, we explore a new paradigm for ensuring fair evaluation—generating adversarial examples which preserve the structure and difficulty of the original questions aimed for assessment, but are unsolvable by LLMs. Focusing on the domain of math word problems, we leverage abstract syntax trees to structurally generate adversarial examples that cause LLMs to produce incorrect answers by simply editing the numeric values in the problems. We conduct experiments on various open- and closed-source LLMs, quantitatively and qualitatively demonstrating that our method significantly degrades their math problem-solving ability. We identify shared vulnerabilities among LLMs and propose a cost-effective approach to attack high-cost models. Additionally, we conduct automatic analysis on math problems and investigate the cause of failure to guide future research on LLM’s mathematical capability.  [^1] Data and code are available: https://github.com/ruoyuxie/adversarial_mwps_generation
LLM-Resistant Math Word Problem Generation via Adversarial Attacks
Roy Xie      Chengxuan Huang      Junlin Wang      Bhuwan Dhingra Duke University {ruoyu.xie, jonathan.huang, junlin.wang2}@duke.edu {bdhingra}@cs.duke.edu
@@ 1 Introduction
At the same time, adversarial attacks on LLMs have gained more attention due to increased awareness of the potential risks associated with LLMs. Most work on adversarial attacks focuses on developing prompts to elicit specific outputs from LLMs (Wang et al., 2021; Zhu et al., 2023). Recent work suggests that even the most powerful aligned LLMs are vulnerable to such attacks (Zou et al., 2023; Carlini et al., 2023). Hence, we might expect that as LLMs become stronger, detecting their outputs will become more difficult, but adversarial examples may still persist Ilyas et al. (2019).
@@
To this end, we introduce a new paradigm for generating homework assignments that LLMs cannot solve by utilizing adversarial attacks. We focus on math word problems (MWPs), which present a unique and challenging intersection of language and mathematics. In this work, we ask the question: “Can LLMs still solve an MWP after changing its numeric values?” We aim to create MWPs that LLMs are unable to solve while maintaining the coherence and original difficulty of the problems. This goal diverges from typical adversarial attacks in machine learning, where small, often imperceptible changes are made to input data (images, text, etc.) to deceive models into making incorrect predictions. Here we do not focus on making imperceptible changes, but rather on preserving the coherence, style, and, most importantly, difficulty of the original problems. Our aim is not just to challenge LLMs but to do so in a way that reflects real-world educational standards to ensure the MWPs remain educationally valuable and relevant.
We generate adversarial examples by editing the numeric values in MWPs. While conceptually simple, doing this effectively is challenging for a couple of reasons: (i) the answer of the resulting problem changes and must be recomputed to assess the robustness of LLMs; and (ii) preserving the difficulty and validity of the modified problem requires us to ensure that all the intermediate calculations in the new problem are consistent with the original problem. For example, if the original problem did not have fractions as intermediate values, in an educational context we would not want the modified problem to have such values either. To tackle these challenges, we first convert MWPs to a code representation and leverage abstract syntax trees (ASTs) to systematically generate adversarial problems.
We evaluate several major LLMs on their math-solving ability under adversarial settings, demonstrating the effectiveness of our method with a significant degradation across various LLMs. We investigate universal attacks and attack transferability, proposing a cost-effective approach to attack high-cost models, achieving a 90% request reduction rate without sacrificing performance. We verify the validity of our generated problems by conducting human evaluations. Lastly, we conduct in-depth automatic analyses to investigate the mathematical vulnerabilities in LLMs and find, unsurprisingly, that the values of the answer and the number of operations have strong correlations with correctness.
@@
Figure 1: Method Overview: given a MWP that an LLM can correctly solve, our method first transforms it into Python code. Next, the Python code is converted into an AST representation which is used to generate adversarial problems by modifying the numeric values in the problem. We place constraints on the nodes of the AST representation to ensure that the modified problem maintains the same difficulty level as the original problem. Despite this, we find that the resulting adversarial examples cause LLMs to predict incorrect answers.
@@ 2 Background and Related Work
The increasing sophistication of LLMs has made it challenging for plagiarism detection tools to accurately identify content generated by machines (Chaka, 2023; Liang et al., 2023). Previous work suggests that detecting LLM-generated content may become nearly impossible as LLMs improve (Sadasivan et al., 2023). The issue extends beyond text-based content; various other disciplines are also threatened by the advancing capabilities of LLMs. For instance, LLMs can now provide step-by-step solutions for MWPs, which complicates the matter of fair evaluation, as students may pass off machine-generated solutions as their own without understanding the underlying mathematical concepts. It remains an open question of how to ensure fair evaluation for students in the presence of LLMs.
@@
The capability of LLMs to solve math problems is currently a hot topic (OpenAI, 2023; Yu et al., 2023; Xu et al., 2023). Studies found that LLMs can significantly improve their math-solving ability through prompt engineering (Yu et al., 2023; Imani et al., 2023), such as chain-of-thought (CoT) (Wei et al., 2022). Converting the MWPs’ solving steps into code or symbolic representations can also improves LLM’s performance (Li et al., 2023; He-Yueya et al., 2023; Gao et al., 2023). While different prompting strategies exist, our focus is to generate adversarial examples that LLMs are unable to solve. Therefore, we aim to minimize prompt engineering by using a unified prompt with zero-shot CoT for all queries in this work. [^2] Prompt template is in Appendix A.2.
@@
Adversarial attacks on LLMs focus on modifying prompts to cause LLMs to output desired content (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023). Similarly, our focus is to modify the MWPs to cause LLMs to output incorrect answers. Bubeck et al. (2023) conducts a limited number of memorization tests on MWPs by randomly changing the numeric values between -10 and 10, suggesting that state-of-the-art LLMs do not simply rely on memorizing problems but on applying a general solution method. However, the number of examples in Bubeck et al. (2023) is relatively small to reach such a conclusion, and we will show in our work that modifying numbers in MWPs does cause LLMs to fail on a larger scale. Zhou et al. (2023), on the other hand, rephrases MWPs by changing the words within the problem and keeping all the numeric values the same. However, this approach makes it challenging to preserve the problem’s original context and structure, and sometimes could introduce inconsistent and problematic content, as shown in Appendix B. Additionally, Zhou et al. (2023) also requires extensive human validation to ensure modification quality. Therefore, we focus on changing numerical values, which ensures the preservation of the underlying logic. To the best of our knowledge, this is the first study to undertake such an endeavor.
@@ 3 Methodology
MWPs are presented using natural language, which creates challenges for systematic structural and syntactic modifications. In this section, we describe our approach to generating adversarial MWPs. An overview of our method can be found in Figure 1.
3.1 Preliminaries
We denote an original problem-answer pair as $p=(x,y)$, where $x$ is a sequence of tokens $x_{1},\dots,x_{n}$ in the problem and $y$ is its ground-truth answer. We denote the set of all possible adversarial modifications to it as:
@@ 3.1 Preliminaries
Here $G$ is a ground-truth function that computes the correct answer for a math problem (in our case this is generated Python code). $F$ is a function which maps the elements of a sequence as: $F(\tilde{x}_{i})\sim\mathbb{R}$ (i.e., a random real number) if $x_{i}$ is numeric and $F(\tilde{x_{i}})=x_{i}$ otherwise. This entire set of adversarial examples will also consist of many unnatural and difficult problems, hence later in §3.3 we will introduce filtering constraints for selecting the adversarial examples that preserve difficulty.
@@ 3.2 Mapping From MWPs to Code to Tree
To structurally modify MWPs, we first convert MWPs into Python code by prompting GPT-4-turbo to generate the code based on the problem-solving steps provided by the datasets. [^3] Prompt can be found in Appendix A.1. As a sanity check, we first select 200 MWPs each from GSM8K (Cobbe et al., 2021) and MultiArith (Roy and Roth, 2015) and check if the generated program results in correct answers. Given the execution steps in code are almost always deterministic, encouragingly, we found only 2 and 5 generated codes have incorrect logic or answers in GSM8K and MultiArith, respectively. We further investigate the generated Python code in §5.3.
Next, we utilize AST, a tree representation for the abstract syntactic structure of programming language code, to convert the generated Python code into a tree. We build the ASTs by traversing through the Python code and constructing nodes corresponding to each statement. The final print statement that outputs the answer is the root of the tree. There are mainly two types of nodes: operation nodes, which carry out the operations in the tree and are the non-leaf nodes, and variable nodes, which correspond to the numeric values from the problem and are the leaf nodes. We discuss the nodes in detail in Appendix C.
@@ 3.3 Adversarial Example Generation
In an adversarial setting, attackers can perform multiple attacks on the system. Similarly, we generate multiple adversarial examples which are different versions of the original MWPs to attack LLMs.
@@
It’s essential to maintain the original difficulty and coherence of the MWPs despite changes in their numeric values. For instance, changing one-digit multiplication to four-digit multiplication significantly alters the problem’s complexity. Furthermore, to keep the mathematical logic intact, the order of magnitude of numbers in the problem should remain unchanged. For example, altering a problem from “subtracting 2 apples from 5” to “subtracting 4 apples from 2” introduces the concept of negative numbers, potentially misaligning with the original problem’s context. Lastly, changes should be contextually coherent. For instance, a generated problem with “the first half of 3 people” would be confusing.
@@
To minimize the difference between the original MWP and its adversarial counterpart while adhering to the educational context, we define a set of Boolean constraints for each node to control the generation quality. Given the original node value $h$ and the new value of node $h^{\prime}$, a newly generated problem is valid if and only if the $h^{\prime}$ satisfies the same set of constraints that $h$ has, for all nodes. A list of node constraints is:
@@
The node constraints only ensure that the adversarial examples are valid in the numerical sense; instances that don’t make logical sense like “$28$-hour work day”, which is not common, can still be generated. Furthermore, different values for the variable nodes can lead to vastly different difficulty levels, as $h^{\prime}$ can be significantly larger than $h$.
To mitigate this, we propose three generation methods to distinguish generated adversarial examples with varying levels of difficulty. We decompose the values $h\approx a\times 10^{b}$, where $a$ is an integer that is not divisible by 10 and with at most $c$ digits. Depending to the numbers $h$, $a$, and $b$, the generation methods are defined as follows:
  • $M1$: Free Generation: regardless of the value of $a$, $a^{\prime}$ is drawn from a uniform distribution between $1$ and $10^{c}$. If the variable node $v$ is a divisor of a division node, then $a^{\prime}$ is drawn from a uniform distribution between $1$ and $\left\lfloor 10^{c/2}\right\rfloor$. The newly generated value $h^{\prime}$ will be $a^{\prime}\times 10^{b}$. The reason for picking smaller numbers for a divisor is that the probability of picking a number from $1$ and $10^{c}$ that is divisible for another number from $1$ and $10^{c}$ is very low for big values of $c$.
  • $M2$: Count of Digits: this generation method ensures that the newly generated value $h^{\prime}$ has the same number of digits as $h$. For example, for $h=100$, some possible $h^{\prime}$ are 942, 589, or 264. This ensures $h^{\prime}$ is always within a reasonable range relative to $h$ and maintains similar problem difficulty levels. For more variability, $h^{\prime}$ can range from 1 to 99 if $h$ has only one digit. For decimal variable values like $h=1.25$, we will use $a=125$ to generate numbers like $a^{\prime}=473$, then convert them back to $h^{\prime}=4.73$.
  • $M3$: Count of Scientific Numbers: this generation method tries to ensure that the new value shares a similar scientific digit count and is within a similar range to the original value. For example, for $h=1500$, $h^{\prime}$ can be 1700, 800, 1200. The rationale is that changing a number to a larger one does not necessarily make calculations harder, but more scientific numbers do. An example would be comparing the numbers 150,000 and 172,568.
$M3$ is the most restrictive generation, followed by $M2$, and $M1$ is the least restrictive, therefore:
@@
The more restrictive a method is, the closer the difficulty levels and coherence are between the adversarial examples and the original problem. Therefore, we use $M3$ as our main generation method as it ensures the adversarial examples comply with the constraints while also gauging the difficulty of problems. Although generations from $M1$ and $M2$ may not fit in the educational context, we are interested in studying LLMs’ math-solving abilities by simply changing the numeric values. We discuss how different methods impact model performance in §4.4. We describe each generation method in more detail in Appendix D and present some sample generations based on each method in Table 5.
@@ 4.1 Experimental Setup
We generate adversarial examples from GSM8K and MultiArith. Both datasets are commonly used for evaluating LLMs’ mathematical capabilities, and MultiArith is sourced from math worksheets used by elementary school students to practice math problems (Roy et al., 2015).
@@
We conduct experiments on 7 open-source models: MetaMath 7B, 70B (Yu et al., 2023), Mistral 7B (Jiang et al., 2023), Llama-2 13B (Touvron et al., 2023), WizardMath 13B (Xu et al., 2023), Vicuna 13B (Chiang et al., 2023), CodeLlama (Roziere et al., 2023), and 2 closed-source models: GPT-4-Turbo (OpenAI, 2023) and GPT-3.5-Turbo (OpenAI, 2022). [^4] Specifically, we use GPT-4-0125-preview and GPT-3.5-turbo-0125. For simplicity, we refer to them as GPT-4 and GPT-3.5 for the rest of the paper. The model selection was largely based on the LLMs’ math-solving ability. Our intention is to evaluate a wide range of LLM performances, including the math-tuned LLMs such as MetaMath and WizardMath, popular open-source LLMs such as Llama-2 and Mistral, and the state-of-the-art API-based GPT models.
@@
Given a set of original problem-answer pairs $P$ and a LLM $L$, we use the follwing evaluation metrics:
  • Original Accuracy (OA): the accuracy of $L$ on the original problems:
@@
  • Attack Accuracy (AA): given an indicator function $I_{xy}$:
    $I_{xy}=\mathrm{1}\left[\forall(\tilde{x},\tilde{y})\in A(x,y):L(\tilde{x})=\tilde{y}\right],$ (4)
    $L$’s accuracy on the adversarial examples:
    $\text{AA}(L)=\frac{\sum_{(x,y)\in P}I_{xy}}{|P|}.$ (5)
  • Attack Success Rate (ASR): the relative decrease in accuracy due to adversarial modifications:
@@
4.2 Rephrasing Attacks
We examine the effectiveness of a commonly used rephrasing approach, where the modifications are centered around the words within the MWPs. Zhou et al. (2023) freeze the logical entities (e.g., entity names and values) and iteratively go through each unfrozen token to swap it with similar tokens that could cause LLMs to fail. They then validate the adversarial example with human manual checking to ensure the example’s quality. Zhou et al. (2023) present the RobustMath dataset, which includes 214 original problems from a combination of GSM8K and MultiArith, and generate 300 rephrased versions. As a baseline, we report their ASR in Table 2. We observe that rephrasing attacks result in lower accuracy for 4 out of the 7 models; however, some models show improved accuracy, suggesting that such attacks may not generalize across different LLMs.
@@ 4.2 Rephrasing Attacks
Table 1: Main Results: we generate adversarial examples based on the given constraint. Overall, simply changing the numeric values in the original math problems causes a consistent performance drop across all models, even with the most restrictive generation method.
4.3 Number of Adversarial Examples
We investigate how the number of adversarial examples impacts model performance. Figure 2 shows the average number of possible adversarial examples per question for each constraint.
Figure 3: Increase Examples: we report the model performance given various number of adversarial examples. For several models, only 10 adversarial examples are enough to degrade the performance.

### v2

@@ Abstract
Large language models (LLMs) have significantly transformed the educational landscape. As current plagiarism detection tools struggle to keep pace with LLMs’ rapid advancements, the educational community faces the challenge of assessing students’ true problem-solving abilities in the presence of LLMs. In this work, we explore a new paradigm for ensuring fair evaluation—generating adversarial examples which preserve the structure and difficulty of the original questions aimed for assessment, but are unsolvable by LLMs. Focusing on the domain of math word problems, we leverage abstract syntax trees to structurally generate adversarial examples that cause LLMs to produce incorrect answers by simply editing the numeric values in the problems. We conduct experiments on various open- and closed-source LLMs, quantitatively and qualitatively demonstrating that our method significantly degrades their math problem-solving ability. We identify shared vulnerabilities among LLMs and propose a cost-effective approach to attack high-cost models. Additionally, we conduct automatic analysis on math problems and investigate the cause of failure, offering a nuanced view into model’s limitation.
@@ 1 Introduction
At the same time, adversarial attacks on LLMs have gained more attention due to increased awareness of the potential risks associated with LLMs. Most work on adversarial attacks focuses on developing prompts to elicit specific outputs from LLMs (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023) Recent work suggests that even the most powerful aligned LLMs are vulnerable to such attacks (Zou et al., 2023; Carlini et al., 2023). Hence, we might expect that as LLMs become stronger, detecting their outputs will become more difficult, but adversarial examples may still persist (Ilyas et al., 2019; Wei et al., 2023).
@@
To this end, we introduce a new paradigm for generating homework assignments that LLMs cannot solve by utilizing adversarial attacks. We focus on math word problems (MWPs), which present a unique and challenging intersection of language and mathematics. In this work, we ask the question: “Can LLMs still solve an MWP after changing its numeric values?” We aim to create MWPs that LLMs are unable to solve while maintaining the original difficulty and coherence of the problems. Our aim is not just to challenge LLMs but to do so in a way that reflects real-world educational standards to ensure the MWPs remain educationally valuable and relevant.
@@
Figure 1: Method Overview Given a MWP that an LLM can correctly solve, our method first transforms it into Python code. The Python code then is converted into an AST representation which is used to generate adversarial problems by modifying the numeric values in the problem. We place constraints on the nodes of the AST to ensure that the modified problem maintains the same difficulty level as the original problem. Despite this, we find that the resulting adversarial examples cause LLMs to predict incorrect answers.
While conceptually simple, doing this automatically is challenging as directly modifying numbers in the problem could lead to nonsensical problems. For example, consider the problem “A class has 6 male students and half as many female students. How many students in total?” Directly change the number of male students from 6 to 5 (or any odd numbers) without checking its intermediate computational steps might result in “2.5 female students,” which is not only illogical but also introducing fraction to the problem, which might change the indented difficulty. Similarly, changing the number from 6 to 624 might make the problem unrealistic and much harder for students. To generate meaningful and coherent problem variations, it is essential to consider the logical implications of number modifications and ensure that the resulting problem remains plausible and solvable. We further discuss the importance of educational constraints in §3.2.
To effectively generate modifications in large scale to assess the robustness of LLMs also requires: (i) The answer to the altered problem changes must be recomputed automatically; and (ii) Preserving the difficulty and validity of the modified problem requires us to ensure that all the intermediate calculations in the new problem are consistent with the original problem. To tackle these challenges, we first convert MWPs to a code representation and leverage abstract syntax trees (ASTs) to map each calculation step into a node. We then define educational constraints for each node to ensure all the desired properties are preserved for the generated new problem.
In this work, we evaluate several major LLMs and demonstrate the effectiveness of our method by achieving a significant attack success rate (ASR). Our approach outperforms the previous rephrasing attack by an average of 60% ASR. We investigate universal attacks and attack transferability, proposing a cost-effective approach to attack high-cost models (e.g., GPT-4) by reducing API request calls by 90% without sacrificing performance. We conduct human evaluations to verify that our generated problems indeed preserve the original coherence and difficulty. Furthermore, we perform a regression analysis and find that our adversarial examples exploit different weaknesses of each model, offering valuable insights and a nuanced view into model’s limitation.
@@ 2 Background and Related Work
As LLMs become more adept at generating human-like text, it becomes increasingly difficult to distinguish between student-and machine-generated content (Chaka, 2023; Liang et al., 2023), which poses significant challenges for educational institutions in ensuring fair evaluation of student work (Yan et al., 2024). This issue also extends beyond traditional written assignments, as LLMs can now provide detailed solutions to complex problems across various disciplines (Abedi et al., 2023). Consequently, educators must develop new strategies to assess student understanding and maintain the integrity of the evaluation process (Liu et al., 2023).
@@
Our work also relates closely to LLMs math reasoning ability (OpenAI, 2023; Yu et al., 2023; Xu et al., 2023). Studies found that LLMs can significantly improve their math-solving ability through prompt engineering (Yu et al., 2023; Imani et al., 2023a), such as chain-of-thought (CoT) (Wei et al., 2022). Converting the MWPs’ solving steps into code or symbolic representations can also improves LLM’s performance (Li et al., 2023; He-Yueya et al., 2023; Gao et al., 2023). [^1] While different prompting strategies exist, our focus is to generate adversarial examples that LLMs are unable to solve. Therefore, we try to minimize prompt engineering by using a unified prompt with zero-shot CoT for all evaluations. Prompt can be found in Appendix A.2.
@@
Adversarial attacks on LLMs involve modifying prompts to alter their behavior (Zhang et al., 2020; Zou et al., 2023; Carlini et al., 2023). Similarly, we modify MWPs to cause LLMs to output incorrect answers. Bubeck et al. (2023) conducted limited memorization tests on MWPs by randomly changing numeric values between -10 and 10, suggesting that state-of-the-art LLMs do not solely rely on memorization but apply general solution methods. However, their study’s sample size was relatively small, and we demonstrate that modifying numbers in MWPs causes LLMs to fail on a larger scale. On the other hand, Zhou et al. (2023) rephrased MWPs by changing words while preserving numeric values. This approach risks altering the original context and introducing inconsistencies and problematic content (see Appendix B), necessitating extensive human validation. Therefore, we focus on changing numerical values, preserving the underlying logic. To the best of our knowledge, this is the first study to undertake such an endeavor.
@@ 3 Methodology
MWPs are presented using natural language, which creates challenges for systematic structural and syntactic modifications. In this section, we describe our approach to generating adversarial MWPs. An overview of our method can be found in Figure 1. We denote an original problem-answer pair as $p=(x,y)$, where $x$ is a sequence of tokens $x_{1},\dots,x_{n}$ in the problem and $y$ is its ground-truth answer. We define $G$ as a ground-truth function that computes the correct answer for a math problem [^2] In our case $G$ is generated Python code. and $F$ as a function which maps the elements of a sequence as: $F(\tilde{x}_{i})\sim\mathbb{R}$ (i.e., a random real number) if $x_{i}$ is numeric and $F(\tilde{x_{i}})=x_{i}$ otherwise. We denote the set of all possible adversarial modifications to it as:
@@
$A(x,y)$ will also consist of many unnatural and difficult problems, hence later in §3.2 we will introduce filtering constraints for selecting the adversarial examples that preserve difficulty.
@@ 3.1 Mapping From MWPs to Code to Tree
To structurally modify MWPs, we first use GPT-4 to generate the Python code that reflects the solution steps given the problem and its final answer. [^3] We only require the problem and its final answer, which is almost always given in the context of math problems. Prompt can be found in Appendix A.1. As a sanity check, we select 200 MWPs each from GSM8K (Cobbe et al., 2021) and MultiArith (Roy & Roth, 2015) and check if the generated program results in correct answers. Given the execution steps in code are almost always deterministic, encouragingly, we found only 1% and 2.5% generated codes have incorrect logic or answers in GSM8K and MultiArith, respectively. We further investigate the generated Python code in §5.
Next, we utilize AST, a tree for the abstract syntactic structure of programming language code, to convert the generated Python code into a tree representation. We build the ASTs by traversing through the Python code and constructing nodes corresponding to each statement. The final print statement that outputs the answer is the root of the tree. Each AST has mainly two types of nodes: operation nodes, which carry out the operations in the tree and are the non-leaf nodes, and variable nodes, which correspond to the numeric values from the problem and are the leaf nodes. We discuss the nodes in AST in detail in Appendix C.
@@ 3.2 Adversarial Example Generation
In an adversarial setting, attackers can perform multiple attacks on the system, and any successful attack signifies the possible flaws of the system (Carlini & Wagner, 2017). Similarly, we generate adversarial examples which are different versions of the original MWPs to attack LLMs.
@@
It’s essential to maintain the original difficulty and coherence of the MWPs despite changes in their numeric values. For instance, changing one-digit multiplication to four-digit multiplication significantly alters the problem’s complexity. To maintain mathematical logic intact, the order of magnitude of numbers in the problem should also remain unchanged. For example, changing a problem from “Jack ate 2 out of a total of 5 apples” to “Jack ate 5 out of a total of 2 apples” not only introduces the concept of negative numbers but also alters the problem’s logical structure and coherency. Furthermore, such modifications can lead to unrealistic scenarios that deviate from the intended learning objectives and may cause confusion.
@@
To minimize the difference between the original MWP and its adversarial counterpart while adhering to the educational context, we define a set of Boolean constraints for each node to control the generation quality. Given the original node value $h$ and its new value $h^{\prime}$, a newly generated problem is valid if and only if $h^{\prime}$ satisfies the same set of constraints that $h$ has, for all nodes. A list of node constraints is:
@@
The node constraints ensure the adversarial examples are valid in the numerical sense; instances that don’t make logical sense like “$28$-hour work day”, which is not common, can still be generated. Different values for the variable nodes can also lead to vastly different difficulty levels, as $h^{\prime}$ can be significantly larger than $h$.
To mitigate this, we propose three generation methods to distinguish generated adversarial examples with different levels of difficulty. We decompose the values $h\approx a\times 10^{b}$, where $a$ is an integer that is not divisible by 10 and with at most $c$ digits. Depending on the numbers $h$, $a$, and $b$, the generation methods are defined as follows:
  • $M1$: Free Generation: Allowing the generated problem to have a wide range of numbers with minimal constraints, with each $h^{\prime}$ to be $a^{\prime}\times 10^{b}$. Regardless of the value of $a$, $a^{\prime}$ is drawn from a uniform distribution between $1$ and $10^{c}$. If the variable node $v$ is a divisor of a division node, then $a^{\prime}$ is drawn from a uniform distribution between $1$ and $\left\lfloor 10^{c/2}\right\rfloor$.
  • $M2$: Count of Digits: Constraining the generated value $h^{\prime}$ has the same number of digits as $h$. For example, for $h=100$, some possible $h^{\prime}$ are 942, 589, or 264. This ensures $h^{\prime}$ is always within a reasonable range relative to $h$ and maintains similar problem difficulty levels. For more variability, $h^{\prime}$ can range from 1 to 99 if $h$ has only one digit. For decimal variable values like $h=1.25$, we use $a=125$ to generate numbers like $a^{\prime}=473$, then convert them back to $h^{\prime}=4.73$.
  • $M3$: Count of Scientific Numbers: Constraining the new value shares a similar scientific digit count and is within a similar range to the original value. For example, for $h=1500$, $h^{\prime}\sim\text{Pois}(h)$ can be 1700, 800, 1200, where Pois is the Poisson distribution. The rationale is that larger numbers do not result in more difficult problems, but more scientific numbers do. An example would be comparing the numbers 150,000 and 172,568.
$M3$ is the most restrictive generation, followed by $M2$, and $M1$ is the least restrictive:
@@
The more restrictive a method is, the closer the difficulty levels and coherence are between the an adversarial example and its original version. Therefore, we use $M3$ as our main generation method as it ensures the adversarial examples comply with all constraints to maintain the original difficulty and coherence. Although generations from $M1$ and $M2$ may not fit in the educational context, we are interested in studying LLMs’ math-solving abilities by simply changing the numeric values. We discuss how different methods impact model performance in §4.2. We describe each generation method in more detail in Appendix D and present the generated examples based on three methods in Table 5.
@@ 4.1 Experimental Setup
We generate problem variants from GSM8K (Cobbe et al., 2021) and MultiArith (Roy & Roth, 2015). Both datasets are commonly used for evaluating LLMs’ mathematical capabilities, and MultiArith is sourced from math worksheets used by elementary school students to practice math problems (Roy et al., 2015). [^4] While MWPs exhibit different levels of difficulty, we focus on MWPs at this particular difficulty level as a proof of concept in this work.
@@
We conduct experiments on 7 open-source models: MetaMath 7B, 70B (Yu et al., 2023), Mistral 7B (Jiang et al., 2023), Llama 2 13B (Touvron et al., 2023), WizardMath 13B (Xu et al., 2023), Vicuna 13B (Chiang et al., 2023), CodeLlama (Roziere et al., 2023), and 2 closed-source models: GPT-4 (OpenAI, 2023) and GPT-3.5 (OpenAI, 2022). [^5] Specifically, we use GPT-4-0125-preview and GPT-3.5-turbo-0125. For simplicity, we refer to them as GPT-4 and GPT-3.5 for the rest of the paper. The model selection was largely based on the LLMs’ math-solving ability. Our intention is to evaluate a wide range of LLM performances, including the math-tuned LLMs such as MetaMath and WizardMath, popular open-source LLMs such as Llama 2 and Mistral, and the API-based GPT models.
@@
We follow the previous adversarial attack literature to measure the attacks, which looks for at least one perturbation that fools the model (Croce et al., 2020; Pruthi et al., 2019; Jia & Liang, 2017). A problem is considered incorrect if it has at least one incorrect variation. Such variation is also known as an adversarial example. Given a set of original problem-answer pairs $P$ and a LLM $L$, we define: [^6] OA, AA, ASR are reported in % in this work.
  • Original Accuracy (OA): the accuracy of $L$ on the original problems,
@@
  • Attack Accuracy (AA): the accuracy of $L$ on the adversarial modifications,
    $\text{AA}(L)=\frac{\sum_{(x,y)\in P}\mathrm{1}\left[\forall(\tilde{x},\tilde{y})\in A(x,y):L(\tilde{x})=\tilde{y}\right]}{|P|}.$ (4)
  • Attack Success Rate (ASR): relative decrease in accuracy due to adversarial modifications,
@@
Table 1: Main Attacks We generate MWPs based on different generation methods. Overall, simply changing the numeric values in the problems causes a consistent performance drop across all LLMs, even with the most restrictive generation method.
Baseline
As a baseline, we examine the effectiveness of a commonly used rephrasing approach, where the modifications are centered around the words within the MWPs. Zhou et al. (2023) freeze the logical entities and iteratively swap each unfrozen token with similar tokens that cause LLMs to fail. They then validate the adversarial example with human manual checking. Zhou et al. (2023) created the RobustMath dataset, which includes 214 original problems from a combination of GSM8K and MultiArith with their 300 rephrased versions. We report the results in Appendix E, observing ASR for 4 out of the 7 models; however, some models even show improved accuracy given these adversarial examples, suggesting that such attacks may not generalize well across different LLMs.
4.2 Our Attacks
Problem Variation Generation